OpenStudy (anonymous):

The area, a, of an ellipse can be determined using the formula a = πxy, where x and y are half the lengths of the largest and smallest diameters of the ellipse. Which is an equivalent equation solved for y? a. y = a · πx b. y = a + (πx) c. y = a – πx d. y = a ÷ (πx)

OpenStudy (anonymous):

@shinalcantara

OpenStudy (shinalcantara):

\[a = \pi xy \] dividing both sides by pi x \[\frac{ a }{ \pi x } = y\]